OpenStudy (anonymous):

please HELP me i will medal and fan.. The table below shows two equations: Equation 1 |3x − 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 0, 1. The solutions to equation 1 are x = −1.3, 2 and equation 2 has no solution. The solutions to equation 1 are x = −1.3, 2 and equation 2 has solutions x = 0, 1.

OpenStudy (anonymous):

I THINK THAT IT IS D?!

OpenStudy (anonymous):

Do you know the steps for solving absolute vale equations?

OpenStudy (anonymous):

No not really

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

you set the absolute value side equal to the other side as a Case #1. and set the absolute value side equal to the other side all times -1 as a Case #2

OpenStudy (anonymous):

also, absolute value can never get to be negative

OpenStudy (anonymous):

the answer has to be the third one

OpenStudy (anonymous):

just by deduction

OpenStudy (anonymous):

you don't even have to solve :)

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

Yes, but why is there no solution? there is always a solution

OpenStudy (anonymous):

\[\left| -3 \right|= 3\] so the \[\left| x \right|= 3\] means that x could be -3 or positive 3

OpenStudy (anonymous):

ok.. thanks!!

OpenStudy (anonymous):

to solve for absolute values you need to write it as a positive value as well as a negative value \[\left| 3x-1 \right|+7 =2\]can be re-written as 3x - 1 + 7 = 2, 3x = -4, x = -4/3 = -1.33 and -(3x - 1) + 7 = 2 = -3x + 1 + 7 =2 -3x +8 = 2 -3x = -6 x = 2 Does this make sense?

OpenStudy (anonymous):

Try following the same steps for the second equation \[\left| 2x + 1 \right| + 4 = 3\] If you post your work I'll help you through it.